OpenStudy (anonymous):

In the month of October, the population of a species of fish in a river increased by a factor of 1.05 every day. The function below shows the number of fish in the river f(x) after x days: f(x) = 100(1.05)x Which of the following is a reasonable domain for the function? 0 ≤ x ≤ 100 0 ≤ x ≤ 31 All positive integers greater than 100 All positive integers less than 100

OpenStudy (anonymous):

@NeedHelpQuick546

OpenStudy (anonymous):

What is the smallest possible value for f(x)? What is the greatest possible value for f(x)?

OpenStudy (anonymous):

@nelsonjedi so b, right?

OpenStudy (anonymous):

because 30 x 1.05 is 31.25

OpenStudy (anonymous):

x has to be a positive number greater than zero because it begins on the first day of October, Thus if x is greater then 0,,,then f(x) has to be what?

OpenStudy (anonymous):

so, yes

OpenStudy (anonymous):

It is not b

OpenStudy (anonymous):

What if a made x= .001..what would f(x) be?

OpenStudy (anonymous):

well it says greater than OR equal to so it could be 0, therefore, it is b

OpenStudy (anonymous):

You are looking at x we need to figure f(x)...So if x = .01 then f(x) = 100(1.05)(.01)

OpenStudy (anonymous):

..... please, do continue, because now i'm a little confused

OpenStudy (anonymous):

Ok x in the function represents the number of days correct.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

f(x) represents the number of fish,

jimthompson5910 (jim_thompson5910):

The function should be \[\Large f(x) = 100(1.05)^x\] and not \[\Large f(x) = 100(1.05)(x)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, from all of this, i am now thinking its either D or A

OpenStudy (anonymous):

We need to determine the domain of f(x).(The number of fish)

OpenStudy (anonymous):

Can the number of days ever be 0?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How can it be 0?

OpenStudy (anonymous):

o days pass

OpenStudy (anonymous):

Look at the equation..how many fish do you think are in the lake when we start?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

Bingo..and is that number going to get bigger or smaller?

OpenStudy (anonymous):

...ok i see your point

OpenStudy (anonymous):

never 0 because it would make the population 0 in the long run of the equation

jimthompson5910 (jim_thompson5910):

that would be true IF the equation were \[\Large f(x) = 100(1.05)(x)\] but the equation is \[\Large f(x) = 100(1.05)^x\]

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

@jim_thompson5910 so it is A?

jimthompson5910 (jim_thompson5910):

hint: it has to do with October

OpenStudy (anonymous):

My apologies I did not read the question correctly I was trying to solve the range...my bad.

OpenStudy (anonymous):

lol its ok

OpenStudy (anonymous):

and yea it was B the entire time lol

OpenStudy (anonymous):

:-}